Consider the following augmented matrix [ 1 -1 2 1]

[3 -1 5 -2]

[-4 2 x^2-8 x+2]


Determine the values of x for which the system has

I) no solution,

Ii) exactly one solution,

Iii) infinitely many solutions


1
Expert's answer
2021-05-24T03:01:03-0400

Augmented matrix 


"A=\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 3 & -1 & 5 & -2 \\
-4 & 2 & x^2-8 & x+2
\end{bmatrix}"



"R_2=R_2-3R_1"


"\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 0 & 2 & -1 & -5 \\
-4 & 2 & x^2-8 & x+2
\end{bmatrix}"

"R_3=R_3+4R_1"


"\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 0 & 2 & -1 & -5 \\
0 & -2 & x^2 & x+6
\end{bmatrix}"

"R_2=R_2\/2"


"\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 0 & 1 & -1/2 & -5/2 \\
0 & -2 & x^2 & x+6
\end{bmatrix}"

"R_1=R_1+R_2"


"\begin{bmatrix}
 1 & 0 & 3/2 & -3/2 \\
 0 & 1 & -1/2 & -5/2 \\
0 & -2 & x^2 & x+6
\end{bmatrix}"

"R_3=R_3+2R_2"


"\begin{bmatrix}
 1 & 0 & 3/2 & -3/2 \\
 0 & 1 & -1/2 & -5/2 \\
0 & 0 & x^2-1 & x+1
\end{bmatrix}"

I) no solution


"\begin{matrix}
 x^2-1=0 \\
 x+1\not=0
\end{matrix}=>x=1"

Ii) exactly one solution


"x^2-1\not=0=>x\not=\pm1"

Iii) infinitely many solutions


"\begin{matrix}
 x^2-1=0 \\
 x+1=0
\end{matrix}=>x=-1"